OpenStudy (hockeychick23):

what is _7C_4 (the 7 and 4 are subscripts)

OpenStudy (campbell_st):

so you are asking for the 7C4 combination...? \[^7C_{4} = \frac{7!}{4!(7 - 3)!}\] just calculate it out

OpenStudy (hockeychick23):

ok so i got 5040 for the top half, but can you tell me what to distribute first to get the bottom half please?

OpenStudy (hockeychick23):

oh and also the 7 is a subscript also not a power

OpenStudy (anonymous):

\[^{7}C_{4} = \frac{ 7! }{ 4!(7-4)! }\] \[^{7}C_{4} = \frac{ 7! }{ 4!*3! }\] \[^{7}C_{4} = \frac{ 7*6*5*4! }{ 4!*3! }\] 4! in denominator will cancel out the 4! in numerator 3! = 3*2*1 = 6 \[^{7}C_{4} = \frac{ 7*6*5 }{ 6 }\] \[^{7}C_{4} = { 7*5 }\]

OpenStudy (hockeychick23):

thank you!!

OpenStudy (anonymous):

anytime :)